# Matchings in hypercubes extend to long cycles

### v1

@@ 4.3.2. $M$ does not contain $z$-dangerous (near) half-layers
We proceed explaining the further arguments assuming that there is a matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ that satisfies the first condition stated above. The arguments in the other case are analogous, but not completely symmetric (due to $z$), so our algorithm has to compute both cases separately. As mentioned before, we can assume that $M$ is maximal in the sense that it covers at least one end vertex of every edge of $Q_{5}\setminus z$, and therefore $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ is maximal in $Q_{1,0}^{\prime}\setminus z$.
@@
We also compute the following sets in a loop for $i=1,2,3,4$.
  • $\mathcal{M}_{i}^{\prime}$: The subset of pairs $(M,A)$ from $\mathcal{M}_{i}$ such that $M$ cannot be extended to a linear forest of $K(Q_{1,0}^{\prime})$ with terminals exactly in $A$.
  • ¯$\overline{\mathcal{M}}_{i}$: The set of matchings $M$ of $K(Q_{5})$ for which there is a pair $(M^{\prime},A^{\prime})$ in $\mathcal{M}_{i}^{\prime}$ such that $M$ induced on $Q_{1,0}^{\prime}$ equals $M^{\prime}$ and the set of end vertices of edges in $M$ that leave the set $Q_{1,0}^{\prime}$ is $A^{\prime}$.
  • ¯$\overline{\mathcal{M}}^{\prime}_{i}$: The matchings in ¯$\overline{\mathcal{M}}_{i}$ that cannot be extended to a cycle $K(Q_{5}\setminus z)$.
  • $\mathcal{M}_{i+1}$: The set of pairs $(M,A)$ obtained from some pair $(M^{\prime},A^{\prime})$ in $\mathcal{M}_{i}^{\prime}$ by removing one of the edges from $M^{\prime}$ and adding the two end vertices of the removed edge to the set $A$.
We note that $\mathcal{M}_{1}=\mathcal{M}_{1}^{\prime}$. Our program reports that the sets ¯$\overline{\mathcal{M}}^{\prime}_{i}$ for $i=1,2,3,4$ are all empty. This proves that $M$ can be extended to a cycle in $K(Q_{5})$ that avoids $z$, along the lines argued in Case 1. The running time for this part of our program is approximately 10 hours.
@@ 4.4.1. Generating matchings
We now explain how our program generates different matchings using the data structure $M$. We build $M$ successively by adding matching edges one after the other. In the simplest case $M$ has no edges initially, but a set $A$ of vertices of even cardinality marked as $\operatorname{\mathtt{MATCH}}$ in $M$. In other cases, some edges may already be prescribed or some vertices may be excluded, which is achieved by marking the corresponding vertices as $\operatorname{\mathtt{TERMINAL}}$, $\operatorname{\mathtt{UNCOVERED}}$, or $\operatorname{\mathtt{FORBIDDEN}}$. We use two approaches for generating all perfect matchings on the set $A$. The first one is based on depth-search first (DFS) and the second one on breadth-search first (BFS). DFS takes $M$, chooses a vertex $u\in A$, and for every other vertex $v\in A$ calls DFS for the matching obtained from $M$ by adding the edge $uv$ (by setting $M[u]:=v$ and $M[v]:=u$). This approach stores only the current (partial) matching in memory, and tests a matching for extendability once it is completed to a perfect matching on $A$. Only the ones that are not extendable are stored for later processing.
@@
Define $\mathcal{S}:=\{M\}$. Let $a$ be the number of $\operatorname{\mathtt{MATCH}}$ vertices in $M$. for $a/2$ times do
       Define $\mathcal{S}^{\prime}:=\emptyset$. foreach $N\in\mathcal{S}$ do
             Let $u$ be a vertex marked as $\operatorname{\mathtt{MATCH}}$ in $N$. foreach vertex $v\neq u$ marked as $\operatorname{\mathtt{MATCH}}$ in $N$ do
                   Let $N^{\prime}$ be the matching obtained from $N$ by adding the edge $uv$. Set $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}\cup\{N^{\prime}\}$.
      Set $\mathcal{S}:=\mathcal{S}^{\prime}$ Remove isomorphic matchings from $\mathcal{S}$.
@@
BFS stores an entire set $\mathcal{S}$ of matchings simultaneously, which is more memory-intensive, but has the advantage that it allows removing isomorphic matchings each time an edge has been added (see the last line in the pseudocode above). We therefore gain running time if we can quickly remove a significant number of isomorphic matchings. Intuitively, the fraction of removed matchings decreases with the number of edges added, i.e., while many matchings with few edges will be removed as isomorphic, in later stages only few matchings with more edges will be removed. For optimal performance, we apply BFS as long there is enough memory available and the isomorphism testing removes a significant number of matchings. Afterwards we continue building the matching via DFS.
@@ 4.4.2. Extending matchings to a cycle or linear forest
       Choose a vertex $u$ such that $M[u]$ is a vertex or $\operatorname{\mathtt{TERMINAL}}$. foreach $i\in[d]$ do
@@
                   Let $M^{\prime}$ be obtained from $M$ by adding the edge $uu^{i}$. if extends($M^{\prime}$) then
@@
Adding an edge $uu^{i}$ is simple, although we have to consider few cases depending on the values of $M[u]$ and $M[u^{i}]$. If both $M[u]$ and $M[u^{i}]$ are vertices such that $M[u]\neq u^{i}$, we join the vertices $M[u]$ and $M[u^{i}]$ by an edge and we mark $u$ and $u^{i}$ as $\operatorname{\mathtt{FORBIDDEN}}$. Observe that the resulting matching $M^{\prime}$ can be extended if and only if $M\cup\{uu^{i}\}$ can be extended. I.e., in our algorithm we store in $M$ a matching that is a compact representation of the actual partial extension, obtained by replacing every path by the shortcut between its two end vertices. Similarly, if $M[u]$ is a vertex and $M[u^{i}]=\operatorname{\mathtt{UNCOVERED}}$, we join vertices $M[u]$ and $u^{i}$ by an edge and we mark $u$ as $\operatorname{\mathtt{FORBIDDEN}}$. If $M[u]$ is a vertex and $M[u^{i}]=\operatorname{\mathtt{TERMINAL}}$, we mark $u$ and $u^{i}$ as $\operatorname{\mathtt{FORBIDDEN}}$ and $M[u]$ as $\operatorname{\mathtt{TERMINAL}}$. The remaining cases are analogous.
It remains to determine whether an edge $uu^{i}$ can be added and whether it is the last edge to be added. Clearly, if $M[u]=u^{i}$, then adding the edge $uu^{i}$ closes a cycle. In order to avoid closing a cycle that misses some edges of $M$, we maintain two counters for the number of edges and terminals in $M$. The last edge $uu^{i}$ is added if $M$ has one edge $uu^{i}$ and no $\operatorname{\mathtt{TERMINAL}}$ vertices (extension to cycle), or $M$ has no edges and two $\operatorname{\mathtt{TERMINAL}}$ vertices $u$ and $u^{i}$ (extension to linear forest). We can add an edge $uu^{i}$ (not the last one) to the matching $M$ if $M[u]\neq u^{i}$ and $M$ has at least four terminals or at least one of $u$ and $M[u]$ is not marked as $\operatorname{\mathtt{TERMINAL}}$.
@@
4.4.3. Removing isomorphic matchings
We follow the approach from [GNŠ18] to remove isomorphic matchings from a given set $\mathcal{S}$ of matchings of $Q_{d}$. It is well-know that every automorphism of $Q_{d}$ is composed of a transposition (i.e., switching certain directions) and a permutation of directions, i.e., there are $2^{d}d!$ automorphisms of $Q_{d}$. However, in our setting we require the fixed vertex $z=\emptyset$ to be forbidden, so we consider only automorphisms that permute directions, sometimes with one or two directions fixed. Our function removes isomorphic matchings from $\mathcal{S}$ by applying to every matching $M\in\mathcal{S}$ all relevant automorphisms and by selecting the lexicographically minimal representation. As isomorphic matchings have the same lexicographically minimal representation, this will remove duplicates.

### v2

@@ 4.3.2. $M$ does not contain $z$-dangerous (near) half-layers
We proceed to explain the further arguments assuming that there is a matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ that satisfies condition (a) stated above. The arguments in case (b) are analogous, but not completely symmetric (due to $z$), so our algorithm has to compute both cases separately. As mentioned before, we can assume that $M$ is maximal in the sense that it covers at least one end vertex of every edge of $Q_{5}\setminus z$, and therefore $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ is maximal in $Q_{1,0}^{\prime}\setminus z$.
@@
We also compute the following sets in a loop for $k=1,2,3,4$.
  • $\mathcal{M}_{k}^{\prime}$: The subset of pairs $(M,A)$ from $\mathcal{M}_{k}$ such that $M$ cannot be extended to a linear forest of $K(Q_{1,0}^{\prime}\setminus z)$ with terminals exactly in $A$.
  • ¯$\overline{\mathcal{M}}_{k}$: The set of matchings $M$ of $K(Q_{5})$ for which there is a pair $(M^{\prime},A^{\prime})$ in $\mathcal{M}_{k}^{\prime}$ such that $M$ induced on $Q_{1,0}^{\prime}$ equals $M^{\prime}$ and the set of end vertices of edges in $M$ that leave the set $Q_{1,0}^{\prime}$ is $A^{\prime}$.
  • ¯$\overline{\mathcal{M}}^{\prime}_{k}$: The matchings in ¯$\overline{\mathcal{M}}_{k}$ that cannot be extended to a cycle of $K(Q_{5}\setminus z)$.
  • $\mathcal{M}_{k+1}$: The set of pairs $(M,A)$ obtained from some pair $(M^{\prime},A^{\prime})$ in $\mathcal{M}_{k}^{\prime}$ by removing one of the edges from $M^{\prime}$ and adding the two end vertices of the removed edge to the set $A$.
We note that $\mathcal{M}_{1}=\mathcal{M}_{1}^{\prime}$. Our program reports that the sets ¯$\overline{\mathcal{M}}^{\prime}_{k}$ for $k=1,2,3,4$ are all empty. This proves that $M$ can be extended to a cycle in $K(Q_{5})$ that avoids $z$, as argued in the following.
Recall the assumption from case (a) that $|A_{1,0}^{\prime}|\geq 2$ is even and there exists a perfect matching $P_{1,0}^{\prime}$ on $K(A_{1,0}^{\prime})$ such that $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ does not contain half-layers of $Q^{i}_{0}$ nor the half-layer $L^{j}_{1i}$ of $Q^{j}_{1}$. This means that $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}\in\mathcal{M}$.
If $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}\notin\mathcal{M}^{\prime}$, then $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ can be extended to a cycle $C^{\prime}$ of $K(Q_{1,0}^{\prime}\setminus z)$. Let $P_{1,0}$ be the shortcut edges of the linear forest $(C^{\prime}\setminus P_{1,0}^{\prime})\cup M_{1,0,-}$ in $Q_{1,0}$. By Theorem 5, $P_{1,0}\cup M_{1,0}$ can be extended to a cycle $D$ of $K(Q_{1,0})$. Therefore, $(C^{\prime}\setminus P_{1,0}^{\prime})\cup M_{1,0,-}\cup(D\setminus P_{1,0})$ is a cycle in $K(Q_{5})$ that extends $M$ and avoids $z$.
If $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}\in\mathcal{M}^{\prime}$, then let $N_{1}$ be a matching that contains all edges of $M_{1,0}^{\prime}\cup P_{1,0}^{\prime}$ except one edge of $P_{1,0}^{\prime}$ and let $A_{1}$ be the end vertices of the removed edge. By definition, we have $(N_{1},A_{1})\in\mathcal{M}_{1}$. If $(N_{1},A_{1})\notin\mathcal{M}_{1}^{\prime}$, then $N_{1}$ be extended to a linear forest of $K(Q_{1,0}^{\prime}\setminus z)$ with terminals exactly in $A_{1}$. Similarly as in the previous paragraph, we then construct shortcut edges in $Q_{1,0}$, use Theorem 5 to obtain a cycle in $K(Q_{1,0})$ which we combine with the linear forest to obtain a cycle in $K(Q_{5})$ that extends $M$ and avoids $z$.
If $(N_{1},A_{1})\in\mathcal{M}_{1}^{\prime}$ and $|M_{1,0,-}|=2$, then we have $M\in\overline{\mathcal{M}}_{1}$ by definition. Since ¯$\overline{\mathcal{M}}^{\prime}_{1}$ is empty, $M$ can be extended to a cycle of $K(Q_{5}\setminus z)$.
If $(N_{1},A_{1})\in\mathcal{M}_{1}^{\prime}$ and $|M_{1,0,-}|\geq 4$, then let $N_{2}$ be a matching that contains all edges of $N_{1}$ except another edge of $P_{1,0}^{\prime}$ and let $A_{2}$ be the end vertices of both edges of $P_{1,0}^{\prime}$ excluded in $N_{2}$. By definition, we have $(N_{2},A_{2})\in\mathcal{M}_{2}$. If $(N_{2},A_{2})\notin\mathcal{M}_{2}^{\prime}$ or $|M_{1,0,-}|=4$, then we proceed similarly to before. Otherwise, we construct and process $N_{3}$ and $A_{3}$ analogously. If $(N_{3},A_{3})\notin\mathcal{M}_{3}^{\prime}$ or $|M_{1,0,-}|=6$, then we proceed similarly to before. Otherwise, we construct and process $N_{4}$ and $A_{4}$ analogously. Since $Q_{1,0}$ has 8 vertices, this is the last possible case, i.e., we have $(N_{4},A_{4})\notin\mathcal{M}_{4}^{\prime}$ or $|M_{1,0,-}|=8$.
The running time for this part of our program is approximately 10 hours for each of the two cases (a) and (b).
@@ 4.4.1. Generating matchings
We now explain how our program generates different matchings using the data structure $M$. We build $M$ successively by adding matching edges one after the other. In the simplest case $M$ has no edges initially, but a set $A$ of vertices of even cardinality marked as $\operatorname{\mathtt{MATCH}}$ in $M$. In other cases, some edges may already be prescribed or some vertices may be excluded, which is achieved by marking the corresponding vertices as $\operatorname{\mathtt{TERMINAL}}$, $\operatorname{\mathtt{UNCOVERED}}$, or $\operatorname{\mathtt{FORBIDDEN}}$. We use two approaches for generating all perfect matchings on the set $A$. The first one is based on depth-first search (DFS) and the second one on breadth-first search (BFS). DFS takes $M$, chooses a vertex $u\in A$, and for every other vertex $v\in A$ calls DFS for the matching obtained from $M$ by adding the edge $uv$ (by setting $M[u]:=v$ and $M[v]:=u$). This approach stores only the current (partial) matching in memory, and tests a matching for extendability once it is completed to a perfect matching on $A$. Only the ones that are not extendable are stored for later processing.
@@
Define $\mathcal{S}:=\{M\}$
Let $a$ be the number of $\operatorname{\mathtt{MATCH}}$ vertices in $M$
for $a/2$ times do
       Define $\mathcal{S}^{\prime}:=\emptyset$
       foreach $N\in\mathcal{S}$ do
             Let $u$ be a vertex marked as $\operatorname{\mathtt{MATCH}}$ in $N$
             foreach vertex $v\neq u$ marked as $\operatorname{\mathtt{MATCH}}$ in $N$ do
                   Let $N^{\prime}$ be the matching obtained from $N$ by adding the edge $uv$
                   Set $\mathcal{S}^{\prime}:=\mathcal{S}^{\prime}\cup\{N^{\prime}\}$
      Set $\mathcal{S}:=\mathcal{S}^{\prime}$
       Remove isomorphic matchings from $\mathcal{S}$
@@
BFS stores an entire set $\mathcal{S}$ of matchings simultaneously, which is more memory-intensive, but has the advantage that it allows removing isomorphic matchings each time an edge has been added (see the last line in the pseudocode above). We therefore gain running time if we can quickly remove a significant number of isomorphic matchings. Intuitively, the fraction of removed matchings decreases with the number of edges added, i.e., while many matchings with few edges will be removed as isomorphic, in later stages only few matchings with more edges will be removed. For optimal performance, we apply BFS as long as there is enough memory available and the isomorphism testing removes a significant number of matchings. Afterwards we continue building the matching via DFS.
We follow the approach from [GNŠ18] to remove isomorphic matchings from a set $\mathcal{S}$ of matchings of $Q_{d}$. It is well-know that every automorphism of $Q_{d}$ is composed of a transposition (i.e., switching certain directions) and a permutation of directions, i.e., there are $2^{d}d!$ automorphisms of $Q_{d}$. However, in our setting we require the fixed vertex $z=\emptyset$ to be forbidden, so we consider only automorphisms that permute directions, sometimes with one or two directions fixed. Our function removes isomorphic matchings from $\mathcal{S}$ by applying to every matching $M\in\mathcal{S}$ all relevant automorphisms and by selecting the lexicographically minimal representation. As isomorphic matchings have the same lexicographically minimal representation, this will remove duplicates.
@@ 4.4.2. Extending matchings to a cycle or linear forest
       Choose a vertex $u$ such that $M[u]$ is a vertex or $\operatorname{\mathtt{TERMINAL}}$
       foreach $i\in[d]$ do
@@
                   Let $M^{\prime}$ be obtained from $M$ by adding the edge $uu^{i}$
                   if extends($M^{\prime}$) then
@@
Adding an edge $uu^{i}$ is simple, although we have to consider a few cases depending on the values of $M[u]$ and $M[u^{i}]$. If both $M[u]$ and $M[u^{i}]$ are vertices such that $M[u]\neq u^{i}$, we join the vertices $M[u]$ and $M[u^{i}]$ by an edge and we mark $u$ and $u^{i}$ as $\operatorname{\mathtt{FORBIDDEN}}$. Observe that the resulting matching $M^{\prime}$ can be extended if and only if $M\cup\{uu^{i}\}$ can be extended. I.e., in our algorithm we store in $M$ a matching that is a compact representation of the actual partial extension, obtained by replacing every path by the shortcut between its two end vertices. Similarly, if $M[u]$ is a vertex and $M[u^{i}]=\operatorname{\mathtt{UNCOVERED}}$, we join vertices $M[u]$ and $u^{i}$ by an edge and we mark $u$ as $\operatorname{\mathtt{FORBIDDEN}}$. If $M[u]$ is a vertex and $M[u^{i}]=\operatorname{\mathtt{TERMINAL}}$, we mark $u$ and $u^{i}$ as $\operatorname{\mathtt{FORBIDDEN}}$ and $M[u]$ as $\operatorname{\mathtt{TERMINAL}}$. The remaining cases are analogous.
It remains to determine whether an edge $uu^{i}$ can be added and whether it is the last edge to be added. Clearly, if $M[u]=u^{i}$ and $uu^{i}$ is not an edge of the original matching (but a shortcut edge), then adding the edge $uu^{i}$ closes a cycle. In order to avoid closing a cycle that misses some edges of $M$, we maintain two counters for the number of edges and terminals in $M$. The last edge $uu^{i}$ is added if $M$ has one edge $uu^{i}$ and no $\operatorname{\mathtt{TERMINAL}}$ vertices (extension to cycle), or $M$ has no edges and two $\operatorname{\mathtt{TERMINAL}}$ vertices $u$ and $u^{i}$ (extension to linear forest). We can add an edge $uu^{i}$ (not the last one) to the matching $M$ if $M[u]\neq u^{i}$ and $M$ has at least four terminals or at least one of $u$ and $M[u]$ is not marked as $\operatorname{\mathtt{TERMINAL}}$.
@@ 5. Proofs of Theorems 6, 7, and 12
6. Open questions
We conclude this paper with some interesting directions for future investigations.
It would be interesting to translate the construction described in this paper to an algorithm that computes a cycle $C$ that extends a given matching $M$ of $K(Q_{d})$ in polynomial time (polynomial in the size of the graph $Q_{d}$ or $K(Q_{d})$), and we do not see any fundamental obstacles to such an endeavor. Furthermore, can this be improved, along the lines described in [Fin20], where part of the output can be computed already while knowing only part of the input?
Our Theorems 6 and 7 provide a lower bound for the length of a cycle $C$ that extends a given matching $M$. Similarly, note that our proof of Theorem 8 starts by the assumption that $M$ is (H)-maximal, i.e., in a first step we always add edges to $M$ so that we can guarantee the existence of a direction $i\in[d]$ such that many edges of $M$ have direction $i$; recall Lemma 23. This has the effect that even for very small matchings $M$, the cycle $C$ might be very long. Can we instead prove a theorem about a cycle $C$ extending a given matching $M$, such that the length of $C$ is relatively short compared to the size of $M$, i.e., such that $|C|\leq f(|M|)$ for some reasonable function $f$? Obviously, we will always have $|C|\geq 2|M|$.
More generally, given a matching $M$ of $Q_{d}$ or $K(Q_{d})$, for which integers $\ell$ is there a cycle $C$ with $|C|=\ell$ that extends $M$? This includes the Ruskey-Savage conjecture as a special case (when $\ell=2^{d}$). This can also be stated as a decision problem: Given $M$ and $\ell$, does $C$ exist, yes or no?
Complementing Theorem 4, it is known that not every matching of $K(Q_{d})$ can be extended to a cycle factor of $Q_{d}$. In fact, Dvořák and Fink [DF19] constructed a matching of $B(Q_{d})$ that cannot be extended to a cycle factor. Is there a matching of $K(Q_{d})$ or $B(Q_{d})$ that is extendable to a cycle factor but not to a Hamilton cycle?